OpenStudy (anonymous):

I am trying to understand how to derive the following two formulae. Would you mind giving me a hint as to how I should go about with doing so? a^3 + b^3 = (a + b)(a^2 – ab + b^2) a^3 – b^3 = (a – b)(a^2 + ab + b^2) I am not sure whether or not it requires you to do long division with the formula ax^3 + bx^2 + cx + d. What I am thinking is that you would take the linear factor (a-b) or (a+b) and divide it by the cubic polynomial equation, but I am not sure as to how I will go about with that.

OpenStudy (amistre64):

it not so much as a deriving that gets you there; its just a matter of after doing so many of them you notice a pattern that emerges which makes the process easier

OpenStudy (amistre64):

a^2 -ab +b^2 ---------------- a+b| a^3 + b^3 -a^3 -a^2b ------------ -a^2b +b^3 +a^2b +ab^2 -------------- +ab^2 +b^3 -ab^2 -b^3 ----------- 0 that would work if need be

OpenStudy (anonymous):

(a+b)^3 = a^3 +3a^2b +3ab^2 + b^3 (Pascal) so a^3 +b^3 = (a+b)^3 -3ab(a+b) = (a+b) ((a+b)^2 -3ab) = (a+b)(a^2 +b^2 -ab)

OpenStudy (anonymous):

what is wrong with just multiplying?

OpenStudy (amistre64):

it tells me the next title is legend

OpenStudy (anonymous):

"what is wrong with just multiplying" Nothing at all, all these things (proofs) depend on where u are allowed to begin. The word "derive" is the clue, tutor code for show something that looks like a proof.

OpenStudy (anonymous):